OpenStudy (anonymous):

6(x-2)-16=3(2x-5)+11

OpenStudy (anonymous):

Distribute 6x-12-16=6x-15+11

OpenStudy (anonymous):

distribute 6x-12-16=6x-15+11 combine x -12-16=-15+11 -28=-4 Not sure if I did something wrong or if you typed it wrong.

OpenStudy (anonymous):

6x-28=6x-26